OpenStudy (anonymous):

an airplane tries to fly due north at 100 m/s but a wind is blowing from the west at 30 m/s. what heading (angle and direction) should the plane take to go due north in spite of the wind?

OpenStudy (anonymous):

since wind is blowing from the west, the direction is due east at 30m/s in this case we can use a vector diagram  so, taking angle as x x = tan^-1 (100/30) = 73.30 deg taking speed as y y = (100^2 +30^2)^0.5 (pythagora's theorem) = 104.4 m/s

OpenStudy (anonymous):

sorry, its